# Topological classification of black holes surrounded by quintessence

### v1

@@ Acknowledgments
The research was partially supported by the National Natural Science Foundation of China (Grant No.12265007) and the Doctoral Foundation of Zunyi Normal University of China (BS [2022] 07, QJJ-[2022]-314). Also, the research of H.H. was supported by the Q-CAYLE project, funded by the European Union-Next Generation UE/MICIU/Plan de Recuperacion, Transformacion y Resiliencia/Junta de Castilla y Leon (PRTRC17.11), and also by project PID2023-148409NB-I00, funded by MICIU/AEI/10.13039/501100011033. Financial support of the Department of Education of the Junta de Castilla y Leon and FEDER Funds is also gratefully acknowledged (Reference: CLU-2023-1-05). Additionally, H. H. is grateful to Excellence project FoS UHK 2203/2025-2026 for the financial support.

### v2

@@ Acknowledgments
We are very grateful to Dr. P. K. Yerra for helpful discussion. We also acknowledge the anonymous referees for their valuable comments on improving our paper. The research was partially supported by the National Natural Science Foundation of China (Grant No.12265007). Also, the research of H.H. was supported by the Q-CAYLE project, funded by the European Union-Next Generation UE/MICIU/Plan de Recuperacion, Transformacion y Resiliencia/Junta de Castilla y Leon (PRTRC17.11), and also by project PID2023-148409NB-I00, funded by MICIU/AEI/10.13039/501100011033. Financial support of the Department of Education of the Junta de Castilla y Leon and FEDER Funds is also gratefully acknowledged (Reference: CLU-2023-1-05). Additionally, H. H. is grateful to Excellence project FoS UHK 2203/2025-2026 for the financial support.